Imagine a share of stock that pays a dividend of $2 at the end of one year, $3 at the end of two years, and then dividends grow at a constant rate of 5% ner year thereafter. If the required return is 10%, we can value this share of stock by finding P2, using D3, then find P0= D1/(1.1) + D/(1.1)2 +P2/(1.1)1 In this formula, it appears as though we ignore all dividends from year three on. Do you agree or disagree and why so? (10 marks)
Answer #1

D1 = $2

D2 = $3

Growth g = 5%

D3 = $3 * 1.05 = $3.15

Rate of return = Ke = 10%

P2 = D3 / (ke - g)

P2 = $3.15 / (10%-5%)

P2 = $63

P0 = D1 / (1.1)^1 + D2 / (1.1)^2 + P2 / (1.1)^2

P0 = 2 / 1.1 + 3 / (1.1)^2 + 63 / (1.1)^2

P0 = 1.82 + 2.48 + 52.07

P0 = $56.36

No, i do not agree that the dividends after year 2 are ignored. P2 is the terminal value of the stock dividends accruing to the investor in perpetuity. It represents the total value of dividends as on the date after applying the terminal growth rate by reducing it from rate of return.